# Convergence of the extended Kalman filter with small and state-dependent noise

### v1

@@ 4 Error estimation for the extended Kalman filter
$-B^{\top}Q^{-1}\Big(g(t,M(t))\ell^{\top}+Q(t)\nabla_{y}h(t,M(t))\Big)^{\top}(\ell\ell^{\top})^{-1}H(t)$
$\leq$ $~\mu_{\varepsilon}(1+|Q(t)|+|Q^{-1}(t)|)|Y(t)-M(t)|^{2},$
because $g$, $\ell$ and $(\ell\ell^{\top})^{-1}$ are bounded.
In addition, $Q(t)+Q^{-1}(t)\preceq Ck(t)I$ implies that $|Q(t)|+|Q^{-1}(t)|\leq C|k(t)|$, therefore we can rewrite the last inequality as
$F^{\top}(t)Q^{-1}(t)F(t)\leq\mu_{\varepsilon}Ck(t)(Y(t)-M(t))^{\top}I(Y(t)-M(t)).$ (47)
And then, from $Q(t)\preceq ck(t)I$, we obtain $Ck(t)Q^{-1}\succeq I$ and the inequality (47) can take the form
$F^{\top}(t)Q^{-1}(t)F(t)\leq\mu_{\varepsilon}Ck^{2}(t)(Y(t)-M(t))^{\top}Q^{-1}(t)(Y(t)-M(t)).$ (48)
The condition (45) is satisfied for some $\mu_{\varepsilon}\longrightarrow 0$. Therefore, the conclusion of lemma 4.4 applies, and we deduce that $Q^{-\frac{1}{2}}(t)(Y(t)-M(t))$ is of order $\sqrt{\varepsilon}$. ∎
Remark 4.5
This remark allows us to use condition (45) instead of condition (41). We use the Cauchy-Schwarz inequality corresponding to the inner product defined by the positive definite matrix $\pazocal{K}^{-1}(t)$.
$\left|\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)F(t)\right|$ $\leq\Big(\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)\Big)^{\frac{1}{2}}\Big(F^{\top}(t)\pazocal{K}^{-1}(t)F(t)\Big)^{\frac{1}{2}}$
$\leq\Big(\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)\Big)^{\frac{1}{2}}\Big(\alpha_{\epsilon}k^{2}(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t)\Big)^{\frac{1}{2}}$
$\leq\alpha^{1/2}_{\epsilon}k(t)\pazocal{X}^{\top}(t)\pazocal{K}^{-1}(t)\pazocal{X}(t).$
Remark 4.6
Note that in the proof above, for the sake of clarity we use $A(t)$ is $(Q(t),k(t))$-stable instead of $A(t)$ is $(\varepsilon Q(t),k(t))$-stable since both are equivalent.
Proof
of Theorem 4.1.
Since we have assumed that $a$ and $(\ell\ell^{\top})^{-1}$ are uniformly elliptic and that $h$ is strongly injective, we have
$Q(t)\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}hQ(t)+a(t,M(t))\succeq c(Q^{2}(t)+I)\succeq c\left(Q(t)+Q^{-1}(t)\right)Q(t).$ (49)
@@
$Q(t)+Q(t)^{-1}\succeq\left(b_{1}p(t)+\dfrac{p^{-1}(t)}{b_{2}}\right)I\succeq b\left(p(t)+p^{-1}(t)\right)I,$ (51)
@@
Let $p(0)$ be the trace of $Q(0)$, as a reminder the dynamics of $Q$ is given by
$\hskip-2.84544pt\dot{Q}(t)=$ $~\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big](t,M(t))Q(t)+Q(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}(t,M(t))$ (52)
$~-Q(t)\Big[\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big](t,M(t))Q(t)+a(t,M(t)).$
@@
$-\operatorname{tr}\left(Q(t)\Big[\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]Q(t)\right)+\operatorname{tr}(a(t,M(t))).$
Since $\nabla_{y}f$, $\nabla_{y}h$, $(\ell\ell^{\top})^{-1}$ and $g$ are bounded, and using the trace inequality (72), we have
$\operatorname{tr}\left(Q(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}(t,M(t))\right)$ $\leq\left|\operatorname{tr}\left(Q(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}(t,M(t))\right)\right|$ (53)
$\leq\left|\left[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\right]^{\top}(t,M(t))\right|\operatorname{tr}\left(Q(t)\right).$ (54)
@@
$\operatorname{tr}\left(Q(t)\left(\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\right)(t,M(t))Q(t)\right)\geq\operatorname{tr}\left(CQ^{2}(t)\right)\geq\dfrac{C}{n}\operatorname{tr}\left(Q(t)\right)^{2},$ (55)
with the last inequality coming from the property (73). Thus, we can write
$\operatorname{tr}\left(Q(t)\left(\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\right)(t,M(t))Q(t)\right)\leq\dfrac{C}{n}\operatorname{tr}\left(Q(t)\right)^{2}.$ (56)
Finally, $a(t,M(t))$ being bounded, we can find three positive constants $C_{1},~C_{2},~C_{3}$ such that
$\operatorname{tr}(\dot{Q}(t))\leq-C_{1}\operatorname{tr}\left(Q(t)\right)^{2}+C_{2}\operatorname{tr}\left(Q(t)\right)+C_{3}.$ (57)
@@
$\dot{p}(t)\leq-C_{1}p^{2}(t)+C_{4}.$ (59)
Then, we can say that the solution of (58) with equality sign will be less than the solution of (59) with the equality sign and both starting at the same initial condition $p(0)$.
Similarly, if $\overline{p}(t)$ denotes the trace of $Q^{-1}(t)$, writing the equation of the dynamic of $Q^{-1}$, and relying on the assumptions that $\nabla_{y}h$ is bounded, $a$ and $(\ell\ell^{\top})^{-1}$ are elliptic, we can find two constants $C^{\prime}_{1}$ and $C^{\prime}_{2}$ such that the trace of $Q^{-1}(t)$ is less than the solution of
$\dot{\overline{p}}(t)=-C^{\prime}_{1}\overline{p}^{2}(t)+C^{\prime}_{2}.$ (60)
To obtain equation (60), we first derive the dynamics of $Q^{-1}$ by taking advantage of the fact that $Q(t)Q^{-1}(t)=I$. This implies that $\dot{Q}(t)Q^{-1}(t)+Q(t)\dot{Q}^{-1}(t)=0$ and thus $\dot{Q}^{-1}(t)=-\left(Q^{-1}\dot{Q}Q^{-1}\right)(t)$.
@@
$\dot{Q}^{-1}(t)=$ $~-Q^{-1}(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big](t,M(t))-\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}(t,M(t))Q^{-1}(t)$
$~+\Big[\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big](t,M(t))-Q^{-1}(t)a(t,M(t))Q^{-1}(t).$
Applying the trace and using the same arguments as above, (60) holds.
@@
$\overline{p}^{-1}(t)I\preceq Q(t)\preceq p(t)I.$ (62)
@@
$\leq$ $~C_{3}(p(t)+\overline{p}(t))(\dfrac{C_{4}}{C_{3}}-p(t)\overline{p}(t))$
@@
for some positive constants $C_{3},C_{4},C_{5}$. Hence, the derivative of $p(t)\overline{p}(t)$ is negative as soon as $p(t)\overline{p}(t)$ is greater than $C_{5}$. Moreover, from the assumption on the eigenvalues of $Q(0)$, $p(0)\overline{p}(0)$ is bounded, thus $p(t)\overline{p}(t)$ is bounded as well. Therefore, the conclusion of Lemma 4.3 holds. ∎
@@ 5 Impact of the initial error on the EKF estimation error
Theorem 5.1
Let $(M(t),Q^{\varepsilon}(t))_{t\geq 0}$ be an extended Kalman filter for the filtering problem (1), with the gain $G$ defined as
$G(t)=\left[g(t)\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\right](\ell\ell^{\top})^{-1}(t,M(t)),$ (63)
@@
  2. (A2)
    functions $f$ and $h$ are $\pazocal{C}^{1}$ with bounded derivatives,
  3. (A3)
@@
    $A(t)=\nabla_{y}f(t,\xi(t))-G(t)\nabla_{y}h(t,\xi(t))$ (64)
    is exponentially stable. More precisely, it satisfies (36) for some $c>0$.
    Fixing some $c_{0}<c$, then it holds
    $|Y(t)-M(t)|_{q}\leq C_{q}\sqrt{\varepsilon}+C_{q}|Y(0)-M(0)|_{q}e^{-c_{0}t}.$ (65)
Remark 5.2
Assumption (A1) and (A2) of this theorem are satisfied for the $SI^{\pm}S$ epidemic model, since we consider constant population sizes and strictly positive compartments sizes. For the third assumption, the expression of $A$ is as follows
$A(\tau)=$ $\varepsilon\beta\left[1-2(X_{1}(0)+\varepsilon\xi(\tau))-(X_{2}(0)+\varepsilon Z(\tau))\right]-(\alpha+\rho^{-})\varepsilon$
$-\alpha^{2}\varepsilon\left(Q(\tau)-(X_{1}(0)+\varepsilon\xi(\tau))\right)\left(\alpha(X_{1}(0)+\varepsilon\xi(\tau))+\rho^{-}(X_{2}(0)+\varepsilon Z(\tau))\right)^{-1}.$
Since $A(\tau)$ is a scalar, it is exponentially stable if it is negative for any $\mathbb{F}^{Z}$-adapted process $\xi$. Hence, the parameters and initial values of the filter must be such that $A(\tau)$ is negative in the estimation time window.
The conclusion of Theorem 5.1 shows that for large $t$, the first term $C_{q}\sqrt{\varepsilon}$ is the dominating part of the upper bound on the right-hand side. This implies that for $t$ large enough the filtering error is uniformly bounded w.r.t. time. The proof of Theorem 5.1 involves the use of the following lemma, which is proven in picard1991efficiency.
Lemma 5.3
@@
    Suppose that there exist a uniformly bounded and elliptic family $\pazocal{K}(t)$, and a constant number $k>0$ such that $A(t)$ is $(\pazocal{K}(t),k)$ stable, then $A(t)$ is exponentially stable and the estimate (36) is satisfied with $c=k/2$.
@@
    Conversely, if $A(t)$ is uniformly bounded and exponentially stable (so that estimate (36) holds for some $c>0$) then for any $k<2c$, there exists a family of uniformly bounded and elliptic processes $\pazocal{K}(t)$ that are adapted to the filtration generated by $A$ and are such that $A(t)$ is $(\pazocal{K}(t),k)$ stable. More generally, if $q_{0}$ is a family of uniformly positive numbers, there exists a family of uniformly elliptic processes $\pazocal{K}(t)$ such that $\pazocal{K}(0)=q_{0}I,$ $A(t)$ is $(\pazocal{K}(t),k)$ stable and
    $\operatorname{tr}(\pazocal{K}(t))\leq C\left(1+q_{0}e^{(-(2c-k)t)}\right).$ (66)
Proof
(of Theorem 5.1) Let $(M(t),Q^{\varepsilon}(t))_{t\geq 0}$ be the extended Kalman filter of our model with gain $G(t)$ defined in the theorem and initial guess $M(0)$. Since $f$ and $h$ are $C^{1}$ with bounded derivatives, there exists a $\mathbb{F}$-adapted process $\xi$ such that
$f(t,Y(t))-f(t,M(t))=\nabla_{y}f(t,\xi(t))\left(Y(t)-M(t)\right),$ (67)
and
$h(t,Y(t))-h(t,M(t))=\nabla_{y}h(t,\xi(t))\left(Y(t)-M(t)\right),$ (68)
@@
$f(t,Y(t))-f(t,M(t))-G(t)\left(h(t,Y(t))-h(t,M(t))\right)=$ $\left(\nabla_{y}f(t,\xi(t))-G(t)\nabla_{y}h(t,\xi(t))\right)\pazocal{X}(t)$
$=$ $A(t)\pazocal{X}(t),$
@@
Therefore, defining $\widetilde{W}(t):=\left(W^{(1)}(t),W^{(2)}(t)\right)^{\top}$, we have
@@
$=\left[f(t,Y(t))-f(t,M(t))-G(t)\left(h(t,Y(t))-h(t,M(t))\right)\right]dt+$
$\phantom{=}~~\sqrt{\varepsilon}\left[\sigma(t,Y(t))dW^{(1)}(t)+\left(g(t,Y(t))-G(t)\ell(t,Y(t))\right)dW^{(2)}(t)\right]$
$\phantom{=}~~\left(\nabla_{y}f(t,\xi_{t})-G(t)\nabla_{y}h(t,\xi(t))\right)\pazocal{X}(t)dt+$
$\phantom{=}~~\sqrt{\varepsilon}\left(\sigma(t,Y(t)),~g(t,Y(t))-G(t)\ell(t,Y(t))\right)d\widetilde{W}(t)$
@@
Since $A(t)$ is exponentially stable and bounded by assumption, Lemma 5.3 asserts that for any $k<2c$, there exists a family of processes $\pazocal{K}(t)$ that are uniformly bounded, elliptic and adapted such that $A(t)$ is $(\pazocal{K}(t),k)$ stable, and for some family of uniformly positive numbers $q_{0}=|Y(0)-M(0)|_{q}^{2}$, $\pazocal{K}(t)$ satisfy
$\operatorname{tr}(\pazocal{K}(t))\leq C\left(1+q_{0}e^{(-(2c-k)t)}\right).$ (69)
We deduce also that $A(t)$ is $(\varepsilon\pazocal{K}(t),k)$ stable. We are then in the setting of Lemma 4.4 with $F(t)=0$. Thus for $\alpha=0$ and because $\sigma,~~g$ and $\ell$ are bounded, for any $q\geq 1$ and $\varepsilon$ small enough, if $\pazocal{X}^{\top}(0)(\varepsilon\pazocal{K}(0))^{-1}\pazocal{X}(0)$ is bounded in $L^{q}$, then the process $\pazocal{X}^{\top}(t)(\varepsilon\pazocal{K}(t))^{-1}\pazocal{X}(t)$ is bounded in $L^{q}$. In addition,
$\varepsilon\pazocal{K}(t)\preceq\operatorname{tr}(\varepsilon\pazocal{K}(t))I$~~ $\text{implies}~~(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1}I\preceq(\varepsilon\pazocal{K}(t))^{-1},$
$\text{thus}~~~\pazocal{X}^{\top}(t)(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1}\pazocal{X}(t)\leq\pazocal{X}^{\top}(t)(\varepsilon\pazocal{K}(t))^{-1}\pazocal{X}(t),$
$\text{then}~~~\pazocal{X}^{\top}(t)(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1}\pazocal{X}(t)\text{ is bounded in }L^{q}.$
And finally, given the fact that $\pazocal{X}^{\top}(t)(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1}\pazocal{X}(t)=|(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1/2}\pazocal{X}(t)|^{2}$, we deduce that $(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1/2}\pazocal{X}(t)$ is bounded in $L^{q}$.
Therefore, there exists $C_{q}$ such that
$|(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1/2}\pazocal{X}(t)|_{q}\leq C_{q},$ (70)

### v2

@@ 4 Error estimation for the extended Kalman filter
$-B^{\top}Q^{-1}\Big(g(t,M(t))\ell^{\top}+Q(t)\nabla_{y}h(t,M(t))\Big)(\ell\ell^{\top})^{-1}H(t)$
The expression of $F^{\top}(t)Q^{-1}(t)F(t)$ involves matrices of the form $Q(\ell\ell^{\top})^{-1}Q^{-1}$ and $Q^{-1}(\ell\ell^{\top})^{-1}Q$. Using the fact that $Q$ and $Q^{-1}$ are symmetric and positive definite, one can show that these two matrices have the same eigenvalues as the matrix $(\ell\ell^{\top})^{-1}$. Since $(\ell\ell^{\top})^{-1}$ is bounded, these matrices are also bounded. Thus, using our assumptions that $g$, $\ell$ and $(\ell\ell^{\top})^{-1}$ are bounded, there exists a family of positive constants $\mu=(\mu^{\varepsilon})_{\varepsilon>0}$ depending on $\overline{\mu}$ such that
$F^{\top}(t)Q^{-1}(t)F(t)\leq$ $~\mu^{\varepsilon}(1+|Q(t)|+|Q^{-1}(t)|)|Y(t)-M(t)|^{2}.$
In addition, $Q(t)+Q^{-1}(t)\preceq Ck(t)I$ implies that $|Q(t)|+|Q^{-1}(t)|\leq 2C|k(t)|$, therefore we can rewrite the last inequality as
$F^{\top}(t)Q^{-1}(t)F(t)\leq 2\mu^{\varepsilon}Ck(t)(Y(t)-M(t))^{\top}I(Y(t)-M(t)).$ (4.16)
Then, from $Q(t)\preceq Ck(t)I$, we obtain $Ck(t)Q^{-1}\succeq I$ and the inequality (4.16) can take the following form (for a new constant C),
$F^{\top}(t)Q^{-1}(t)F(t)\leq\mu^{\varepsilon}Ck^{2}(t)(Y(t)-M(t))^{\top}Q^{-1}(t)(Y(t)-M(t)).$ (4.17)
The condition (4.15) is satisfied for some $\mu^{\varepsilon}\to 0$, which completes the proof of condition (4.4).
Verfication of condition (4.5)
This can be deduced from the assumption (4.2) as follows. We have that $Q(t)+Q^{-1}(t)\preceq Ck(t)I$. This implies that  $D^{\top}(t)(Q(t)+Q^{-1}(t))D(t)\preceq D^{\top}(t)Ck(t)ID(t)$  and thus
$~~D^{\top}(t)Q(t)D(t)+D^{\top}(t)Q^{-1}(t)D(t)$ $\preceq Ck(t)|D(t)|^{2}.$ (4.18)
Since $Q(t)$ and $Q^{-1}(t)$ are positive definite, it holds $D^{\top}(t)Q(t)D(t)\succeq 0$ and also $D^{\top}(t)Q^{-1}(t)D(t)\succeq 0$. In addition, $|D(t)|$ is bounded, thus $D^{\top}(t)Q^{-1}(t)D(t)=O(k(t))$ which proves (4.5).
Final conclusion
In view of Definition 3.12one can see that if $A$ is $(Q,k)$-stable then it is also $(\varepsilon Q,k)$-stable for any $\varepsilon>0$. Thus, we can apply Lemma 4.4 with $\pazocal{K}=\varepsilon Q$ (instead of $\pazocal{K}=Q$) and deduce that $\pazocal{X}^{\top}\pazocal{K}^{-1}\pazocal{X}=(Y(t)-M(t))^{\top}(\varepsilon Q)^{-1}(Y(t)-M(t))$ is bounded in $L^{2}$, hence $\varepsilon^{-1/2}|Q^{-1/2}(t)(Y(t)-M(t))|^{2}\leq C$, for some $C>0$. This implies that $Q^{-1/2}(t)(Y(t)-M(t))$ is of order $\sqrt{\varepsilon}$.
Proof.
(of Theorem 4.1)
Since we have assumed (A4) that $\Phi$} and $(\ell\ell^{\top})^{-1}$ are uniformly elliptic and that $h$ is strongly injective, we have
$Q(t)\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}hQ(t)+{\Phi}(t,M(t))\succeq c(Q^{2}(t)+I)\succeq c\left(Q(t)+Q^{-1}(t)\right)Q(t).$ (4.19)
@@
$Q(t)+Q(t)^{-1}\succeq\left(b_{1}p(t)+\dfrac{p^{-1}(t)}{b_{2}}\right)I\succeq b\left(p(t)+p^{-1}(t)\right)I,$ (4.21)
@@
Let $p(0)$ be the trace of $Q(0)$, as a reminder the dynamics of $Q$ is given by, see (3.11),
$\dot{Q}(t)=-Q(t)\Big[\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big](t,M(t))Q(t)$
$\hskip 10.00002pt+\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big](t,M(t))Q(t)+Q(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}(t,M(t))$
$\hskip 10.00002pt+{\Phi}(t,M(t)).$
@@
$-\operatorname{tr}\left(Q(t)\Big[\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]Q(t)\right)+\operatorname{tr}({\Phi}(t,M(t))).$
Since $\nabla_{y}f$, $\nabla_{y}h$, $(\ell\ell^{\top})^{-1}$ and $g$ are bounded, and using the trace inequality (D.2), we have
$\operatorname{tr}\big(Q(t)\big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\big]^{\top}(t,M(t))\big)\hskip 140.00021pt$ (4.22)
$\leq\big|\operatorname{tr}\big(Q(t)\big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\big]^{\top}(t,M(t))\big)\big|\hskip 3.00003pt$ (4.23)
$\leq\big|\big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\big]^{\top}(t,M(t))\big|\operatorname{tr}\big(Q(t)\big).$ (4.24)
@@
$\operatorname{tr}\left(Q(t)\left(\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\right)(t,M(t))Q(t)\right)\geq\operatorname{tr}\left(CQ^{2}(t)\right)\geq\dfrac{C}{n}\operatorname{tr}\left(Q(t)\right)^{2},$ (4.25)
with the last inequality coming from the property (D.3). Thus, we can write
$\operatorname{tr}\left(Q(t)\left(\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\right)(t,M(t))Q(t)\right)\leq\dfrac{C}{n}\operatorname{tr}\left(Q(t)\right)^{2}.$ (4.26)
Finally, ${\Phi}(t,M(t))$ being bounded, we can find three positive constants $C_{1},~C_{2},~C_{3}$ such that
$\operatorname{tr}(\dot{Q}(t))\leq-C_{1}\operatorname{tr}\left(Q(t)\right)^{2}+C_{2}\operatorname{tr}\left(Q(t)\right)+C_{3}.$ (4.27)
@@
$\dot{p}(t)\leq-C_{1}p^{2}(t)+C_{4}.$ (4.29)
Then, the solution of (4.28) with equality sign will be less than the solution of (4.29) with equality sign and both starting at the same initial condition $p(0)$.
Similarly, if $\overline{p}(t)$ denotes the trace of $Q^{-1}(t)$, writing the equation of the dynamic of $Q^{-1}$, and relying on the assumptions that $\nabla_{y}h$ is bounded, $\Phi$} and $(\ell\ell^{\top})^{-1}$ are elliptic, we can find two constants $C^{\prime}_{1}$ and $C^{\prime}_{2}$ such that the trace of $Q^{-1}(t)$ is less than the solution of
$\dot{\overline{p}}(t)=-C^{\prime}_{1}\overline{p}^{2}(t)+C^{\prime}_{2}.$ (4.30)
To obtain equation (4.30), we first derive the dynamics of $Q^{-1}$ by taking advantage of the fact that $Q(t)Q^{-1}(t)=I$. This implies that $\dot{Q}(t)Q^{-1}(t)+Q(t)\dot{Q}^{-1}(t)=0$ and thus $\dot{Q}^{-1}(t)=-\left(Q^{-1}\dot{Q}Q^{-1}\right)(t)$.
@@
$\dot{Q}^{-1}(t)=$ $-Q^{-1}(t){\Phi}(t,M(t))Q^{-1}(t)-Q^{-1}(t)\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big](t,M(t))$
$-\Big[\nabla_{y}f-g\ell^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big]^{\top}(t,M(t))Q^{-1}(t)+\Big[\nabla_{y}h^{\top}(\ell\ell^{\top})^{-1}\nabla_{y}h\Big](t,M(t)).$
Applying the trace and using the same arguments as above, (4.30) holds.
@@
$\overline{p}^{-1}(t)I\preceq Q(t)\preceq p(t)I.$ (4.32)
@@
$\leq$ $~C_{3}(p(t)+\overline{p}(t))\Big(\dfrac{C_{4}}{C_{3}}-p(t)\overline{p}(t)\Big)$
@@
for some positive constants $C_{3},C_{4},C_{5}$. Hence, the derivative of $p(t)\overline{p}(t)$ is negative as soon as $p(t)\overline{p}(t)$ is greater than $C_{5}$. Moreover, from the assumption (A4) on the eigenvalues of $Q(0)$, $p(0)\overline{p}(0)$ is bounded, thus $p(t)\overline{p}(t)$ is bounded as well. Therefore, the conclusion of Lemma 4.3 holds. ∎
@@ 5 Impact of the initial error on the EKF estimation error
Theorem 5.1.
Let $(M(t),Q^{\varepsilon}(t))_{t\geq 0}$ be an extended Kalman filter for the filtering problem (1.1), with the gain $G$ defined as
$G(t)=\left[g(t)\ell^{\top}(t,M(t))+Q(t)\nabla_{y}h^{\top}(t,M(t))\right](\ell\ell^{\top})^{-1}(t,M(t)),$ (5.1)
@@
  2. (B2)
    the functions $f$ and $h$ are $\pazocal{C}^{1}$ with bounded derivatives,
  3. (B3)
@@
            $A(t)=\nabla_{y}f(t,\xi(t))-G(t)\nabla_{y}h(t,\xi(t))$
    is exponentially stable. More precisely, it satisfies (3.16) for some $c>0$.
Then there exists a constant $C_{q}>0$ such that for all $c_{0}<c$ it holds
$|Y(t)-M(t)|_{q}\leq C_{q}\sqrt{\varepsilon}+C_{q}|Y(0)-M(0)|_{q}e^{-c_{0}t}.$ (5.2)
The conclusion of Theorem 5.1 shows that for large $t$, the first term $C_{q}\sqrt{\varepsilon}$ is the dominating part of the upper bound on the right-hand side. This implies that for $t$ large enough the filtering error is uniformly bounded w.r.t. time. The proof of Theorem 5.1 involves the use of the following lemma, which is proven in [22].
Lemma 5.2.
@@
    Suppose that there exist a uniformly bounded and elliptic family $\pazocal{K}$, and a constant number $k>0$ such that $A$ is $(\pazocal{K},k)$-stable, then $A$ is exponentially stable and the estimate (3.16) is satisfied with $c=k/2$.
@@
    Conversely, if $A$ is uniformly bounded and exponentially stable (so that estimate (3.16) holds for some $c>0$) then for any $k<2c$, there exists a family of uniformly bounded and elliptic processes $\pazocal{K}$ that are adapted to the filtration generated by $A$ and are such that $A$ is $(\pazocal{K},k)$-stable. More generally, if $q_{0}$ is a family of uniformly positive numbers, there exists a family of uniformly elliptic processes $\pazocal{K}$ such that $\pazocal{K}(0)=q_{0}I,$ AA is $(\pazocal{K},k)$-stable and for $t>0$
    $\operatorname{tr}(\pazocal{K}(t))\leq C\left(1+q_{0}e^{(-(2c-k)t)}\right).$ (5.3)
Proof.
(of Theorem 5.1) Let $(M(t),Q^{\varepsilon}(t))_{t\geq 0}$ be the extended Kalman filter of the model with gain $G(t)$ defined in the theorem and initial guess $M(0)$. Since $f$ and $h$ are $C^{1}$ with bounded derivatives by Assumption (B2), there exists a $\mathbb{F}$-adapted process $\xi$ such that
$f(t,Y(t))-f(t,M(t))$ $=\nabla_{y}f(t,\xi(t))\left(Y(t)-M(t)\right),\quad\text{and}$ (5.4)
$h(t,Y(t))-h(t,M(t))$ $=\nabla_{y}h(t,\xi(t))\left(Y(t)-M(t)\right).$ (5.5)
@@
$f(t,Y(t))-f(t,M(t))$ $-G(t)\left(h(t,Y(t))-h(t,M(t))\right)$
$=\left(\nabla_{y}f(t,\xi(t))-G(t)\nabla_{y}h(t,\xi(t))\right)\pazocal{X}(t)=A(t)\pazocal{X}(t),$
@@
Defining $\widetilde{W}(t):=\left(W^{(1)}(t),W^{(2)}(t)\right)^{\top}$, we have
@@
$=$ $\left[f(t,Y(t))-f(t,M(t))-G(t)\left(h(t,Y(t))-h(t,M(t))\right)\right]dt$
$+\sqrt{\varepsilon}\left[\sigma(t,Y(t))dW^{(1)}(t)+\left(g(t,Y(t))-G(t)\ell(t,Y(t))\right)dW^{(2)}(t)\right]$
$=$ $\left(\nabla_{y}f(t,\xi_{t})-G(t)\nabla_{y}h(t,\xi(t))\right)\pazocal{X}(t)dt$
$+\sqrt{\varepsilon}\left(\sigma(t,Y(t)),~g(t,Y(t))-G(t)\ell(t,Y(t))\right)d\widetilde{W}(t)$
@@
Since $A$ is exponentially stable and bounded by Assumption (B3), Lemma 5.2 asserts that for any $k<2c$, there exists a family of processes $\pazocal{K}$ that are uniformly bounded, elliptic and adapted such that $A$ is $(\pazocal{K},k)$ stable, and for some family of uniformly positive numbers $q_{0}=|Y(0)-M(0)|_{q}^{2}$, $\pazocal{K}(t)$ satisfy
$\operatorname{tr}(\pazocal{K}(t))\leq C\big(1+q_{0}e^{-(2c-k)t}\big).$ (5.6)
We deduce also that $A$ is $(\varepsilon\pazocal{K},k)$ stable. We are then in the setting of Lemma 4.4 with $F(t)=0$. Thus for $\eta=0$ and because $\sigma,g$ and $\ell$ are bounded by Assumption (B1), for any $q\geq 1$ and $\varepsilon$ small enough, if $\pazocal{X}^{\top}(0)(\varepsilon\pazocal{K}(0))^{-1}\pazocal{X}(0)$ is bounded in $L^{q}$, then the process $\pazocal{X}^{\top}(t)(\varepsilon\pazocal{K}(t))^{-1}\pazocal{X}(t)$ is bounded in $L^{q}$. In addition, $\varepsilon\pazocal{K}(t)\preceq\operatorname{tr}(\varepsilon\pazocal{K}(t))I$ implies $(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1}I\preceq(\varepsilon\pazocal{K}(t))^{-1}$, thus
$\pazocal{X}^{\top}(t)(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1}\pazocal{X}(t)\leq\pazocal{X}^{\top}(t)(\varepsilon\pazocal{K}(t))^{-1}\pazocal{X}(t).$
It follows that $x^{\top}(t)(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1}\pazocal{X}(t)$ is bounded in $L^{q}$. Finally, given the fact that
$\pazocal{X}^{\top}(t)(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1}\pazocal{X}(t)=|(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1/2}\pazocal{X}(t)|^{2}$, we deduce that $(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1/2}\pazocal{X}(t)$ is bounded in $L^{q}$. Therefore, there exists $C_{q}$ such that
$|(\operatorname{tr}(\varepsilon\pazocal{K}(t)))^{-1/2}\pazocal{X}(t)|_{q}\leq C_{q},$ (5.7)